OpenStudy (anonymous):

Find one (1) solution to the equation 3x2 - 25x + 8 = 0 Round your answer to 2 decimal places, if necessary.

jhonyy9 (jhonyy9):

3x^2 -25x +8=0 - do you know how you calcule the discriminant ? D=b^2 -4ac so can you continue ?